Please help me in my homework

Hillyard Company, an office supplies specialty store, prepares its master budget on a quarterly basis. The following data have been assembled to assist in preparing the master budget for the first quarter:

 

  1. As of December 31 (the end of the prior quarter), the company’s general ledger showed the following account balances:

 



Debits
Credits
Cash$

64,000



Accounts receivable

219,200



Inventory

61,350



Buildings and equipment (net)

374,000



Accounts payable

$

92,325

Common stock


500,000

Retained earnings


126,225


$

718,550

$

718,550


 

  1. Actual sales for December and budgeted sales for the next four months are as follows:

 




December (actual)$

274,000

January$

409,000

February$

606,000

March$

321,000

April$

217,000


 

  1. Sales are 20% for cash and 80% on credit. All payments on credit sales are collected in the month following sale. The accounts receivable at December 31 are a result of December credit sales.

  2. The company’s gross margin is 40% of sales. (In other words, cost of goods sold is 60% of sales.)

  3. Monthly expenses are budgeted as follows: salaries and wages, $39,000 per month: advertising, $57,000 per month; shipping, 5% of sales; other expenses, 3% of sales. Depreciation, including depreciation on new assets acquired during the quarter, will be $45,940 for the quarter.

  4. Each month’s ending inventory should equal 25% of the following month’s cost of goods sold.

  5. One-half of a month’s inventory purchases is paid for in the month of purchase; the other half is paid in the following month.

  6. During February, the company will purchase a new copy machine for $3,400 cash. During March, other equipment will be purchased for cash at a cost of $82,000.

  7. During January, the company will declare and pay $45,000 in cash dividends.

  8. Management wants to maintain a minimum cash balance of $30,000. The company has an agreement with a local bank that allows the company to borrow in increments of $1,000 at the beginning of each month. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. The company would, as far as it is able, repay the loan plus accumulated interest at the end of the quarter.

 

Required:

Using the data above, complete the following statements and schedules for the first quarter:

 

1. Schedule of expected cash collections:

2-a. Merchandise purchases budget:

2-b. Schedule of expected cash disbursements for merchandise purchases:

3. Cash budget:

4. Prepare an absorption costing income statement for the quarter ending March 31.

5. Prepare a balance sheet as of March 31.
Answer #1

(1)

Schedule of Expected Cash Collections




JanuaryFebruaryMarchQuarter
Cash Sales$    79,400$   1,18,800$    61,600$   2,59,800
Credit Sales$ 2,09,600$   3,17,600$ 4,75,200$ 10,02,400
Total Collections$ 2,89,000$   4,36,400$ 5,36,800$ 12,62,200

(2a)

Merchandise Purchases Budget




JanuaryFebruaryMarchQuarter
Budgeted Cost of goods sold$ 2,38,200$   3,56,400$ 1,84,800$   7,79,400
Add desired ending inventory$    89,100$      46,200$    30,750$      30,750
Total needs$ 3,27,300$   4,02,600$ 2,15,550$   8,10,150
Less beginning inventory$    59,550$      89,100$    46,200$      59,550
Required purchases$ 2,67,750$   3,13,500$ 1,69,350$   7,50,600

(2b)

Schedule of Expected Cash Disbursement




JanuaryFebruaryMarchQuarter
December Purchases$    88,725

$      88,725
January Purchases$ 1,33,875$   1,33,875
$   2,67,750
February Purchases
$   1,56,750$ 1,56,750$   3,13,500
March Purchases

$    84,675$      84,675
Total Cash disbursements for purhcases$ 2,22,600$   2,90,625$ 2,41,425$   7,54,650

(3)

Cash Budget




JanuaryFebruaryMarchQuarter
Beginning Cash Balance$    52,000$      30,640$    32,695$      52,000
Add Cash Collections$ 2,89,000$   4,36,400$ 5,36,800$ 12,62,200
Total cash available$ 3,41,000$   4,67,040$ 5,69,495$ 13,14,200
Less cash disbursements:



Purchases of inventory$ 2,22,600$   2,90,625$ 2,41,425$   7,54,650
Selling and administrative expenses$ 1,25,760$   1,41,520$ 1,18,640$   3,85,920
Purchases of equipment
$        2,200$    76,000$      78,200
Cash dividends$    45,000

$      45,000
Total cash disbursements$ 3,93,360$   4,34,345$ 4,36,065$ 12,63,770
Excess (deficiency) of cash$ -52,360$      32,695$ 1,33,430$      50,430
Financing:



Borrowings$    83,00000$      83,000
Repayments00$ -83,000$     -83,000
Interest00$     -2,490$       -2,490
Total financing$    83,0000$ -85,490$       -2,490
Ending cash balance$    30,640$      32,695$    47,940$      47,940

Explanations:

Selling and administrative expenses




JanuaryFebruaryMarchQuarter
Salaries & Wages$    27,000$      27,000$    27,000$      81,000
Advertising$    67,000$      67,000$    67,000$   2,01,000
Shipping$    19,850$      29,700$    15,400$      64,950
Other expenses$    11,910$      17,820$      9,240$      38,970

$ 1,25,760$   1,41,520$ 1,18,640$   3,85,920

(4)

Income Statement

For the Quarter Ended March 31

Sales
$ 12,99,000
Cost of goods sold:

Beginning Inventory$    59,550
Purchases$ 7,50,600
Goods available for sale$ 8,10,150
Ending inventory$    30,750$   7,79,400
Gross margin
$   5,19,600
Selling and administrative expenses:

Salaries and wages$    81,000
Advertising$ 2,01,000
Shipping$    64,950
Depreciation$    44,020
Other expenses$    38,970








$   4,29,940
Net operating income
$      89,660
Interest expense
$        2,490
Net income
$      87,170

(5)

Balance Sheet

March 31

Assets

Current Assets:

Cash
$      47,940
Accounts receivable ($308000 * 80%)
$   2,46,400
Inventory
$      30,750






Total Current assets
$   3,25,090
Building & Equipment, net ($362000 + $2200 + $76000 - $44020)
$   3,96,180
Total assets
$   7,21,270
Liabilities & Stockholders' Equity

Current Liabilities:

Accounts Payable
$      84,675



Stockholders' Equity:

Common Stock$ 5,00,000
Retained Earnings ($94425 + $87170 - $45000)$ 1,36,595





$   6,36,595
Total liabilities & stockholders' equity
$   7,21,270

с D E F January =397000* 20% 209600 =SUM(C3:04) February =594000* 20% =397000*80% =SUM(D3:04) March =308000* 20% =594000*80%=C49 =D49 =E49 76000 2200 45000 =SUM(C29:C32) =C27-C33 =SUM(C30:E30) =SUM(C31:E31) =SUM(C32:E32) =SUM(F29:F32) =F27-F33 =SUM(=D53-D58 Gross margin Selling and administrative expenses: Salaries and wages Advertising Shipping Depreciation Other expenseCurrent Liabilities: Accounts Payable 84675 86 87 88 89 90 Stockholders Equity: Common Stock Retained Earnings ($94425 + $87


answered by: ANURANJAN SARSAM